# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 9.2. The extended Whitehead graph $\widetilde{W}_{v}$ of a virtually free rigid cluster factor $G(v)$
Under the notation of Subsection 9.1, each conjugacy class $\xi\in{\mathcal{H}}_{v}$ is represented by a group $G(u)$ for some $u\in N_{v}$ (see Lemma 9.1). Denote by $n_{\xi}$ the order of the vertex $u$ in the tree $T^{r}_{G}$. This number $n_{\xi}$ is independent of the choice of $G(u)$ as above. Recall also (e.g. from the description in Subsections 3.1 and 3.2) that for each $\xi\in{\mathcal{H}}_{v}$ we have $n_{\xi}\geq 2$.
@@
Let $W_{v}$ be the Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Subsection 9.1. The extended Whitehead graph $\widetilde{W}_{v}$ for ${\mathcal{Q}}_{v}$ is the graph obtained from the Whitehead graph $W_{v}$ as follows. As the vertex set $V_{\widetilde{W}_{v}}$ we take the set $V_{W_{v}}$ of vertices of $W_{v}$. To describe the edge set, consider the labelling $\lambda$ of edges of $W_{v}$, as described in Subsection 8.3, right after Corollary 8.25, and denote it by $\lambda_{v}$ (in order to distinguish it from the analogous labellings for Whitehead graphs $W_{v^{\prime}}$ related to other rigid cluster vertices $v^{\prime}$ of $T^{r}_{G}$). For any $\xi\in{\mathcal{H}}_{v}$, replace each edge $e$ of $W_{v}$ labelled with $\xi$ (i.e. such that $\lambda_{v}(e)=\xi$) by $n_{\xi}-1$ edges $e_{1},\dots,e_{n_{\xi}-1}$ having the same endpoints as $e$. We will call $e_{1},\dots,e_{n_{\xi}-1}$ as above the edges of $\widetilde{W}_{v}$ corresponding to an edge $e$ of $W_{v}$. As part of its structure, equip the so described graph $\widetilde{W}_{v}$ with its own involution system $\tilde{\mathbf{a}}_{v}=(\tilde{a},\{\tilde{\alpha}_{x}\})$ obtained from the corresponding involution system ${\mathbf{a}}_{v}=(a,\{\alpha_{x}\})$ for $W_{v}$ as follows. Identifying naturally the vertices of $\widetilde{W}_{v}$ with the vertices of $W_{v}$, put $\tilde{a}:=a$. Moreover, referring to the notation concerning vertex links as introduced right before Observation 8.26, for any pair of edges $e,e^{\prime}$ of $W_{v}$ such that $\alpha_{x}(e_{x})=e^{\prime}_{a(x)}$, put $\tilde{\alpha}_{x}((e_{i})_{x}):=(e^{\prime}_{i})_{a(x)}$ for $i=1,\dots,n_{\xi}-1$ (where $\xi=\lambda_{v}(e)$, and at the same time $\xi=\lambda_{v}(e^{\prime})$, due to invariance of the labelling under the maps $\alpha_{x}$). Finally, as the last part of the structure of the extended Whitehead graph, we equip it with the labelling $\tilde{\lambda}_{v}$ of the edge set $E_{\widetilde{W}_{v}}$, induced from the labelling $\lambda_{v}:E_{W_{v}}\to{\mathcal{H}}_{v}$ as follows. As the set of values for this new labelling we take the set ~$\widetilde{\mathcal{H}}_{v}:=\{(\xi,j):\xi\in{\mathcal{H}}_{v},1\leq j\leq n_{\xi}-1\}$, and we put $\tilde{\lambda}_{v}(e_{i}):=(\lambda_{v}(e),i)$ for any edge $e$ of $W_{v}$ and any $i=1,\dots,n_{\xi}-1$. Obviously, this induced labelling $\tilde{\lambda}_{v}$ is invariant under the maps $\tilde{\alpha}_{x}$ of the involution system $\tilde{\mathbf{a}}_{v}$, in the same sense as $\lambda_{v}$ is invariant under the maps $\alpha_{x}$. Moreover, again similarly as $\lambda_{v}$, this induced labelling $\tilde{\lambda}_{v}$ associates the same label to some two edges of $\widetilde{W}_{v}$ if and only if the copies of these two edges are involved in the same arm of the $V$-tree ${\mathcal{X}}_{V}((\widetilde{W}_{v},\tilde{\mathbf{a}}_{v}))$.
Next observation is a fairly direct consequence of Proposition 9.3.
@@
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditsch’s JSJ splitting for $G$, and let ${\mathcal{V}}_{v}=(W_{v},{\mathbf{a}}_{v})$ and $\widetilde{\mathcal{V}}_{v}=(\widetilde{W}_{v},\tilde{\mathbf{a}}_{v})$ be the associated connecting $V$-systems.
@@
    The $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ can be obtained from the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{v})$ by leaving the core unchanged and by replacing each arm $A$ in the latter with $n_{\xi}-1$ copies of this arm (sharing only their endpoints, which coincide with the enpoints of $A$), where $\xi=\lambda_{v}(e)$ for any edge $e$ of $W_{v}$ involved in the arm $A$. In particular, the $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ has a similar description to that of ${\mathcal{X}}_{V}({\mathcal{V}}_{v})$, as a tree of arms and the core, as presented in Lemma 8.19, with obvious modification concerning the set of arms.
@@
Since the vertex sets of the graphs $W_{v}$ and $\widetilde{W}_{v}$ canonically coincide, and since the involutions $a:V_{W_{v}}\to V_{W_{v}}$ and $\tilde{a}_{v}:V_{\widetilde{W}_{v}}\to V_{\widetilde{W}_{v}}$ also coincide, there is a canonical identification of the underlying trees $T$ and $\widetilde{T}$ of the tree systems $\Theta[{\mathcal{V}}_{v}]$ and $\Theta[\widetilde{\mathcal{V}}_{v}]$ associated to the connecting $V$-systems ${\mathcal{V}}_{v}$ and $\widetilde{\mathcal{V}}_{v}$. This identification results from the fact that the vertices in both trees correspond to (punctured) copies of the graphs, and edges correspond to gluings of the $V$-peripherals associated to vertices, according to the schemes provided by the $V$-involutions. Under such identification of $T$ with $\widetilde{T}$, it is not hard to observe that to each line $(e_{n}^{\circ})_{n}$ in $\Theta[{\mathcal{V}}_{v}]$ there correspond $n_{\xi}-1$ lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ in the tree system $\Theta[\widetilde{\mathcal{V}}_{v}]$, and that there are no other lines in the latter tree system. Moreover, the geodesics in the tree $\widetilde{T}$ corresponding to the lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ all coincide, and they are equal to the geodesic corresponding to the line $(e_{n}^{\circ})_{n}$ under our identification of $T$ with $\widetilde{T}$. Since the cores under our interest coincide with the Cantor sets appearing as boundaries of the corresponding trees, and since the peripherals in the cores correspond to doubletons given by ends of geodesics induced by lines, part (1) of the lemma clearly follows. Part (2) follows similarly due to the fact that arms in the $V$-trees are induced by lines. ∎
We now pass to discussing some properties of the decomposition of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ into core and arms. We begin with some natural concept of types of the peripherals in the core $C(\widetilde{\mathcal{V}}_{v})$.
@@
Let $P$ be a peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$, and let ${\mathcal{H}}_{v}$ be the set of conjugacy classes of the peripheral subgroups of the abstract factor $G(v)$. Under the identification of $C(\widetilde{\mathcal{V}}_{v})$ with $C({\mathcal{V}}_{v})$ (as given in Lemma 9.6(1)), the type of $P$ is defined as the element $\xi=\lambda_{v}(e)\in{\mathcal{H}}_{v}$, where $e$ is any edge of $W_{v}$ whose copy appears in the line $L\in{\mathcal{L}}_{P}$ (all such edges have the same label).
This assignment of types to the peripherals of the core $C(\widetilde{\mathcal{V}}_{v})$ is natural in the sense provided by the following observation, which follows directly from the definitions and descriptions of the involved objects.
@@
Let $P$ be any peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$, and let $\xi\in{\mathcal{H}}_{v}$ denote its type. Then there is precisely $n_{\xi}-1$ arms in ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ which are attached to $C(\widetilde{\mathcal{V}}_{v})$ along $P$. Moreover, among these arms we have precisely one arm with any label $(\xi,j):1\leq j\leq n_{\xi}-1$ (for the labelling $\tilde{\lambda}_{v}$).
As a final element in this section, we make a record of few further rather obvious observations concerning orientation phenomena in the extended Whitehead graphs and in their associated V-tree systems.
@@
    Note that the edges $e_{1},\dots,e_{n_{\xi}-1}$ which correspond to an edge $e$ of $W_{v}$ in the above description of the extended Whitehead graph $\widetilde{W}_{v}$ share the endpoints with this edge $e$ (under canonical identification of the vertex sets of $W_{v}$ and $\widetilde{W}_{v}$). It follows that any orientation for $e$ induces naturally orientations for all $e_{1},\dots,e_{n_{\xi}-1}$.
@@
    Let $P$ be a peripheral of the core $C(\widetilde{\mathcal{V}}_{v})$. Then it is also a peripheral of the core $C({\mathcal{V}}_{v})$. Recall that $E_{W_{v}}^{P}$ and $E_{\widetilde{W}_{v}}^{P}$ denote then the sets of edges associated to $P$ in the corresponding graphs, as described in Remark/Definition 7.16(2). Using the notation as in the previous remark, we have that $e\in E_{W_{v}}^{P}$ if and only if all edges $e_{1},\dots,e_{n_{\xi}-1}$ belong to $E_{\widetilde{W}_{v}}^{P}$.
@@
    Under notation as in the above remark (2), if we fix an order of the peripheral $P$, we get order-induced orientations at all edges of $\widetilde{W}_{v}$ associated to $P$. Moreover, these orientations are compatible (in the sense of remark (1) above) with the order-induced orientations of the edges of $W_{v}$ associated to $P$ (with respect to the same fixed order of $P$).
10. Structure of a regular tree of graphs for $\partial G$ in terms of the reduced Bowditch’s JSJ splitting of $G$
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose it satisfies condition (3) of Theorem A (i.e. suppose that each rigid cluster factor of the reduced Bowditch JSJ splitting of $G$ is virtually free). In this section we associate to $G$ some graphical connecting system ${\mathcal{R}}_{G}=(\Gamma,{\mathbf{a}},{\mathcal{A}})$ for which, as we will show in Section 13, the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$ is homeomorphic to the Gromov boundary $\partial G$. This connecting system ${\mathcal{R}}_{G}$ will be explicitely determined by various features of the reduced Bowditch JSJ splitting of $G$, and its description constitutes a natural complement to condition (2) of Theorem A, and to condition (2) of Theorem B. We thus rephrase the implication (3$\Rightarrow$(2) in Theorem A as a more precise statement, presented below as Theorem 10.1. (We postpone the proof of Theorem 10.1, and hence of the corresponding implication (3$\Rightarrow$(2) in Theorem A, until Section 13.) We start with describing the graph $\Gamma$ appearing as part of the data in ${\mathcal{R}}_{G}$.
@@
    white vertices of type (v2) which are not adjacent to any rigid cluster vertex.
@@
We now describe the $V$-involution $\mathbf{a}$ appearing as part of data in ${\mathcal{R}}_{G}$. Recall that the extended Whitehead graphs $\widetilde{W}_{z}$ come equipped with $V$-involutions ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Equip also all the $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}\subset\Gamma$ with their standard $V$-involutions (with which they form standard connecting $V$-systems of $\theta$-graphs), as described in Example 7.11, and denote these involutions also by ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Define the $V$-involution ${\mathbf{a}}=(a,\{\alpha_{v}\})$ for $\Gamma$ as follows. The involution $a:V_{\Gamma}\to V_{\Gamma}$ is given by $a(v):=\tilde{a}_{z}(v)$ for each $v\in V_{\Gamma_{z}}\subset V_{\Gamma}$. Moreover, for each $v\in V_{\Gamma}$, if $v$ is a vertex of a component $\Gamma_{z}$, put $\alpha_{v}=(\tilde{\alpha}_{z})_{v}$.

### v2

@@ 9.2. The extended Whitehead graph $\widetilde{W}_{v}$ of a virtually free rigid cluster factor $G(v)$
Under the notation of Section 9.1, each conjugacy class $\xi\in{\mathcal{H}}_{v}$ is represented by a group $G(u)$ for some $u\in N_{v}$ (see Lemma 9.1). Denote by $n_{\xi}$ the degree of the vertex $u$ in the tree $T^{r}_{G}$. This number $n_{\xi}$ is independent of the choice of $G(u)$ as above. Recall also (e.g. from the description in Sections 3.1 and 3.2) that for each $\xi\in{\mathcal{H}}_{v}$ we have $n_{\xi}\geq 2$.
@@
Let $W_{v}$ be the Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Section 9.1. The extended Whitehead graph $\widetilde{W}_{v}$ for ${\mathcal{Q}}_{v}$ is the graph obtained from the $W_{v}$ as follows. As the vertex set $V_{\widetilde{W}_{v}}$ we take the set $V_{W_{v}}$ of vertices of $W_{v}$. To describe the edge set, consider the labelling $\lambda$ of edges of $W_{v}$, as described in Section 8.3, right after Corollary 8.25, and denote it by $\lambda_{v}$ (in order to distinguish it from the analogous labellings for Whitehead graphs $W_{v^{\prime}}$ related to other rigid cluster vertices $v^{\prime}$ of $T^{r}_{G}$). For any $\xi\in{\mathcal{H}}_{v}$, replace each edge $e$ of $W_{v}$ labelled with $\xi$ (i.e. such that $\lambda_{v}(e)=\xi$) by $n_{\xi}-1$ edges $e_{1},\dots,e_{n_{\xi}-1}$ having the same endpoints as $e$. We will call $e_{1},\dots,e_{n_{\xi}-1}$ as above the edges of $\widetilde{W}_{v}$ corresponding to an edge $e$ of $W_{v}$. (As we will see later, the reason for choosing the number $n_{\xi}-1$ comes from the fact that the vertex $u$ of $T^{r}_{G}$, as above, has precisely $n_{\xi}-1$ flexible neighbours, in addition to the single rigid neighbour $v$.) As part of its structure, equip the so described graph $\widetilde{W}_{v}$ with its own involution system $\tilde{\mathbf{a}}_{v}=(\tilde{a},\{\tilde{\alpha}_{x}\})$ obtained from the corresponding involution system ${\mathbf{a}}_{v}=(a,\{\alpha_{x}\})$ for $W_{v}$ as follows. Identifying naturally the vertices of $\widetilde{W}_{v}$ with the vertices of $W_{v}$, put $\tilde{a}:=a$. Moreover, referring to the notation concerning vertex links as introduced right before Observation 8.26, for any pair of edges $e,e^{\prime}$ of $W_{v}$ such that $\alpha_{x}(e_{x})=e^{\prime}_{a(x)}$, put $\tilde{\alpha}_{x}((e_{i})_{x}):=(e^{\prime}_{i})_{a(x)}$ for $i=1,\dots,n_{\xi}-1$ (where $\xi=\lambda_{v}(e)$, and at the same time $\xi=\lambda_{v}(e^{\prime})$, due to invariance of the labelling under the maps $\alpha_{x}$). Finally, as the last part of the structure of the extended Whitehead graph, we equip it with the labelling $\tilde{\lambda}_{v}$ of the edge set $E_{\widetilde{W}_{v}}$, induced from the labelling $\lambda_{v}:E_{W_{v}}\to{\mathcal{H}}_{v}$ as follows. As the set of values for this new labelling we take the set ~$\widetilde{\mathcal{H}}_{v}:=\{(\xi,j):\xi\in{\mathcal{H}}_{v},1\leq j\leq n_{\xi}-1\}$, and we put $\tilde{\lambda}_{v}(e_{i}):=(\lambda_{v}(e),i)$ for any edge $e$ of $W_{v}$ and any $i=1,\dots,n_{\xi}-1$. Obviously, this induced labelling $\tilde{\lambda}_{v}$ is invariant under the maps $\tilde{\alpha}_{x}$ of the involution system $\tilde{\mathbf{a}}_{v}$, in the same sense as $\lambda_{v}$ is invariant under the maps $\alpha_{x}$. Moreover, again similarly as with $\lambda_{v}$, this induced labelling $\tilde{\lambda}_{v}$ associates the same label to some two edges of $\widetilde{W}_{v}$ if and only if copies of these two edges are involved in a common arm of the $V$-tree ${\mathcal{X}}_{V}((\widetilde{W}_{v},\tilde{\mathbf{a}}_{v}))$. In particular, we may say that $\tilde{\lambda}_{v}$ uniquely associates the labels from ~$\widetilde{\mathcal{H}}_{v}$ to arms in ${\mathcal{X}}_{V}((\widetilde{W}_{v},\tilde{\mathbf{a}}_{v}))$.
The next observation is a fairly direct consequence of Proposition 9.3.
@@
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditch JSJ splitting for $G$, and let ${\mathcal{V}}_{v}=(W_{v},{\mathbf{a}}_{v})$ and $\widetilde{\mathcal{V}}_{v}=(\widetilde{W}_{v},\tilde{\mathbf{a}}_{v})$ be the associated connecting $V$-systems.
@@
    The $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ can be obtained from the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{v})$ by leaving the core unchanged and by replacing each arm $A$ in the latter with $n_{\xi}-1$ copies of this arm (sharing only their endpoints, which coincide with the endpoints of $A$), where $\xi=\lambda_{v}(e)$ for any edge $e$ of $W_{v}$ involved in the arm $A$. In particular, the $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ has a description as a tree of arms and the core, $\Psi[\widetilde{\mathcal{V}}_{v}]$, as presented in Lemma 7.12 and in the paragraphs right before it.
@@
Since the vertex sets of the graphs $W_{v}$ and $\widetilde{W}_{v}$ canonically coincide, and since the involutions $a:V_{W_{v}}\to V_{W_{v}}$ and $\tilde{a}_{v}:V_{\widetilde{W}_{v}}\to V_{\widetilde{W}_{v}}$ also coincide, there is a canonical identification of the underlying trees $T$ and $\widetilde{T}$ of the tree systems $\Theta[{\mathcal{V}}_{v}]$ and $\Theta[\widetilde{\mathcal{V}}_{v}]$ associated to the connecting $V$-systems ${\mathcal{V}}_{v}$ and $\widetilde{\mathcal{V}}_{v}$. This identification results from the fact that the vertices in both trees correspond to (punctured) copies of the graphs, and edges correspond to gluings of the $V$-peripherals associated to vertices, according to the schemes provided by the $V$-involutions. Now, let $L=(e_{n}^{\circ})_{n}$ be a line in $\Theta[\widetilde{\mathcal{V}}_{v}]$, and recall that $\xi:=\lambda_{v}(e_{n})$ is then independent of $n$. Under the above identification of $T$ with $\widetilde{T}$, it is not hard to observe that to each line $(e_{n}^{\circ})_{n}$ in $\Theta[{\mathcal{V}}_{v}]$ there correspond $n_{\xi}-1$ lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ in the tree system $\Theta[\widetilde{\mathcal{V}}_{v}]$, and that there are no other lines in the latter tree system. Moreover, the geodesics in the tree $\widetilde{T}$ corresponding to the lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ all coincide, and they are equal to the geodesic corresponding to the line $(e_{n}^{\circ})_{n}$ under our identification of $T$ with $\widetilde{T}$. Since the cores under our interest coincide with the Cantor sets appearing as boundaries of the corresponding trees, and since the peripherals in the cores correspond to doubletons given by ends of geodesics induced by lines, part (1) of the lemma clearly follows. Part (2) follows similarly due to the fact that arms in the $V$-trees are induced by lines. ∎
We now pass to discussing some properties of the decomposition of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ into core and arms. We begin with a natural concept of types of the peripherals in the core $C(\widetilde{\mathcal{V}}_{v})$.
@@
Under notation as in Definition 7.8, let $P=\partial\gamma_{L}$ be a peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$ (which is identified with the core $C(\mathcal{V}_{v})$ as in Lemma 9.6(1)), and let ${\mathcal{H}}_{v}$ be the set of conjugacy classes of the peripheral subgroups of the abstract factor $G(v)$. Recall that ${\mathcal{L}}_{P}$ is the set of lines of $\mathcal{V}_{v}$ whose corresponding arms contain $P$ and that there is necessarily only one such line so that ${\mathcal{L}}_{P}$ is a singleton (see Observation 8.26). The type of $P$ is defined as the element $\xi=\lambda_{v}(e)\in{\mathcal{H}}_{v}$, where $e$ is any edge of $W_{v}$ whose copy appears in the unique line $L\in{\mathcal{L}}_{P}$ (all such edges have the same label, and the line is unique by Observation 8.26(1)).
This assignment of types to the peripherals of the core $C(\widetilde{\mathcal{V}}_{v})$ is natural in the sense provided by the following observation, which can be deduced directly from the definitions and descriptions of the involved objects.
@@
Let $P$ be any peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$, and let $\xi\in{\mathcal{H}}_{v}$ denote its type. Then there is precisely $n_{\xi}-1$ arms in ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ which are attached to $C(\widetilde{\mathcal{V}}_{v})$ along $P$. Moreover, among these arms we have precisely one arm for each label $(\xi,j):1\leq j\leq n_{\xi}-1$ (for the labelling $\tilde{\lambda}_{v}$).
As a final element in this section, we make a record of a few further rather obvious observations concerning orientation phenomena in the extended Whitehead graphs and in their associated V-tree systems.
@@
    Note that the edges $e_{1},\dots,e_{n_{\xi}-1}$ which correspond to an edge $e$ of $W_{v}$ in the above description of the extended Whitehead graph $\widetilde{W}_{v}$ have the same endpoints as $e$ (under the canonical identification of the vertex sets of $W_{v}$ and $\widetilde{W}_{v}$). It follows that any orientation for $e$ induces naturally orientations for all $e_{1},\dots,e_{n_{\xi}-1}$.
@@
    Let $P$ be a peripheral of the core $C(\widetilde{\mathcal{V}}_{v})$. Then it is also a peripheral of the core $C({\mathcal{V}}_{v})$. Recall that $E_{W_{v}}^{P}$ and $E_{\widetilde{W}_{v}}^{P}$ denote then the sets of edges associated to $P$ in the corresponding graphs $W_{v}$ and $\widetilde{W}_{v}$, as described in Remark/Definition 7.16(2). Using the same notation as in the previous remark, we have that $e\in E_{W_{v}}^{P}$ if and only if all edges $e_{1},\dots,e_{n_{\xi}-1}$ belong to $E_{\widetilde{W}_{v}}^{P}$.
@@
    Under notation as in the above remark (2), if we fix an order of the peripheral $P$, we get order-induced orientations at all edges of $\widetilde{W}_{v}$ associated to $P$ (see Observation 8.26(3)). Moreover, these orientations are compatible (in the sense of remark (1) above) with the order-induced orientations of the edges of $W_{v}$ associated to $P$ (with respect to the same fixed order of $P$).
10. Structure of a regular tree of graphs for $\partial G$ in terms of the reduced Bowditch JSJ splitting of $G$
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose it satisfies condition (3) of Theorem A (i.e. suppose that each rigid cluster factor of the reduced Bowditch JSJ splitting of $G$ is virtually free). In this section we associate to $G$ some graphical connecting system ${\mathcal{R}}_{G}=(\Gamma,{\mathbf{a}},{\mathcal{A}})$ (see Definition 6.1) for which, as we will show in Section 13, the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$ is homeomorphic to the Gromov boundary $\partial G$. This connecting system ${\mathcal{R}}_{G}$ will be explicitly determined by various features of the reduced Bowditch JSJ splitting of $G$, and its description constitutes a natural complement to condition (2) of Theorem A, and to condition (2) of Theorem B. We thus rephrase the implication (3$\Rightarrow$(2) in Theorem A as a more precise statement, presented below as Theorem 10.1. (We postpone the proof of Theorem 10.1, and hence of the corresponding implication (3$\Rightarrow$(2) in Theorem A, until Section 13.) We start by describing the graph $\Gamma$ appearing as part of the data in ${\mathcal{R}}_{G}$.
@@
    white vertices of type (v2) (as described in Sections 3.1 and 3.2) which are not adjacent to any rigid cluster vertex.
@@
We now describe the $V$-involution $\mathbf{a}$ appearing as part of data in ${\mathcal{R}}_{G}$. Recall that the extended Whitehead graphs $\widetilde{W}_{z}$ come equipped with $V$-involutions ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Equip also all the $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}\subset\Gamma$ with their standard $V$-involutions (with which they form standard connecting $V$-systems of $\theta$-graphs), as described in the last paragraph of Example 7.11, and denote these involutions also by ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Define the $V$-involution ${\mathbf{a}}=(a,\{\alpha_{v}\})$ for $\Gamma$ as follows. The involution $a:V_{\Gamma}\to V_{\Gamma}$ is given by $a(v):=\tilde{a}_{z}(v)$ for each $v\in V_{\Gamma_{z}}\subset V_{\Gamma}$. Moreover, for each $v\in V_{\Gamma}$, if $v$ is a vertex of a component $\Gamma_{z}$, put $\alpha_{v}=(\tilde{\alpha}_{z})_{v}$.